OpenStudy (keenan):

abs of -3x-2<4 two answers

OpenStudy (anonymous):

\[\left| -3x-2 \right| < 4\]

OpenStudy (keenan):

im sorry its suppose to be a -3 and not -2

OpenStudy (anonymous):

if you were to graph the equation -3x-2 = y and then graph y=4 you could then visualize the area of the graph that must pertain to answer

OpenStudy (anonymous):

\left| -3x-3 \right| < 4

OpenStudy (keenan):

yes

OpenStudy (anonymous):

\[\left| -3x-3 \right| < 4\]

OpenStudy (anonymous):

OK, cool

OpenStudy (keenan):

yep can anyone help me

OpenStudy (anonymous):

sorry, was just graphing

OpenStudy (anonymous):

so basically you to find the value for x such that when you plug into the equation you will get values between -4 and 4, never equal to 4

OpenStudy (anonymous):

let's solve for -3x-3 = 4

OpenStudy (keenan):

ok

OpenStudy (anonymous):

because of the negative slope on the X variable, i.e. (-3). X will have a negative value because a negative multiplied by a negative yields a positive

OpenStudy (anonymous):

so we want to isolate x, therefore, first add 3 to both sides --> \[-3x - 3 + 3 = 4 + 3\]

OpenStudy (anonymous):

--> \[-3x = 7\]

OpenStudy (keenan):

can i stop you for one second ... its suppose to be like a and/or answer

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

we are finding the bounds, the values that x can be such that the equation is equals 4 or -4

OpenStudy (keenan):

like x=4/3 or x>9/3 like that

OpenStudy (anonymous):

so we if found that x = -7/3, we then find that with -3x-3=-4, that x =-1/3

OpenStudy (anonymous):

so between the values \[(-7/3, -1/3)\]

OpenStudy (keenan):

is that suppose to be less then or equal to ...or... greaterthen or equal too

OpenStudy (anonymous):

the equation |-3x-3| will be less than 4

OpenStudy (keenan):

ok

OpenStudy (anonymous):

just check the concept, if x = -2/3, the equations is (-3)*(-2/3) - 3= 2-3 = -1 ... which is less than 4

OpenStudy (anonymous):

so x < -1/3 and x > -7/3

OpenStudy (anonymous):

the notation with the parentheses is another way to represent that expression above

OpenStudy (anonymous):

\[x = (-7/3, -1/3)\] where the parentheses mean that x cannot equal either of those numbers, but it can equal any number between them

OpenStudy (anonymous):

let me know if you need more clarification ;)

OpenStudy (keenan):

it was rong but i posted another problem

OpenStudy (anonymous):

Oh, the answers were incorrect?

OpenStudy (keenan):

they need to be in the correct which on is les then the other

OpenStudy (anonymous):

sorry, looks like I solved the last part incorrectly, should be 1/3

OpenStudy (anonymous):

x > -7/3 and x < 1/3

OpenStudy (anonymous):

so -7/3 < x < 1/3

OpenStudy (keenan):

yeah

OpenStudy (anonymous):

sorry about that